OpenStudy (anonymous):

if f(x) = 3x-2 and g(x)=x^5 find f[g(x)+3]

OpenStudy (hba):

g(x)=x^5 then f(g(x)or f(x^5) will be = ? @needhelp07

OpenStudy (anonymous):

f(x^5)+3???

OpenStudy (anonymous):

@hba

OpenStudy (hba):

Yeah so first find f(x^5) for me We know that f(x)=3x-2

OpenStudy (anonymous):

5x^4????

OpenStudy (hba):

Replace x from x^5

OpenStudy (anonymous):

3(x)^5-2

OpenStudy (hba):

Now add 3

OpenStudy (anonymous):

3(x)^5-2+3 =3(x)^5+1 ????????? @hba

OpenStudy (hba):

Ok you are right intelligent.

OpenStudy (anonymous):

that is the right answer??

OpenStudy (hba):

Yeah you nailed it.

OpenStudy (anonymous):

thx a lot . now i know to solve it

OpenStudy (hba):

Your'e a quick learner :)

OpenStudy (anonymous):

haha tnx again .. need to close my question